OpenStudy (anonymous):

If A and B are independent events, P(A) = 0.4, and P(B) = 0.55, find the probabilities below. P(A union B)_________

OpenStudy (xapproachesinfinity):

P(A or B)=P(A)+P(B)-P(A and B) since there are independent event P(A and B)=0 So P(A or B)=P(A)+P(B) just put in the number and you get the probab you are looking for

OpenStudy (xapproachesinfinity):

P(a and b) is the same as P(a union b), just for clarification

OpenStudy (xapproachesinfinity):

Does this make sense to you?

OpenStudy (anonymous):

no, honestly

OpenStudy (anonymous):

i got it. add the two numbers together then minus the intersection

OpenStudy (kirbykirby):

P(A and B) = P(A)P(B) if they are independent P(A and B) = 0 if A and B are mutually exclusive

OpenStudy (xapproachesinfinity):

Okay let's go slowly then lol P(A union B) means the probability that A occurs or B occurs there is a chance that A will occur or chance of b occurring or may be the two event will both occur what i'm saying now is translated to P(A union B)=P(A)+P(B)-P(A and B)

OpenStudy (xapproachesinfinity):

oh right! P(a and b) i got it wrong case of independent it is as @kirbykirby

OpenStudy (xapproachesinfinity):

thanks for the correction^^

OpenStudy (kirbykirby):

=]

OpenStudy (xapproachesinfinity):

I mixed those two, I haven't practiced this for some time

OpenStudy (xapproachesinfinity):

So our probabilty is P(A or B) = P(A)+P(B)-P(A)P(B) since A and B are independent event P(A and B)= P(A)P(B) like kirbykirby said above

OpenStudy (xapproachesinfinity):

Are you following, or perhaps I confused you lol